Margaret Morris just correctly computed a standard deviation of 6.0 from a sample of 50 observations. What is the variance of Margaret’s sample? : Margaret Morris just correctly computed a standard deviation of 6.0 from a sample of 50 observations. What is the variance of Margaret’s sample?
Answer by oscargut(667)

variance= (standard deviation)^2=6^2=36